OpenStudy (anonymous):

if the numerator of a fraction is increased by 1 , the value of the resulting fraction is 2/3. if the numerator and the denominator are both increased by 5, he value of the resulting fraction is 7/10. find the original fraction.

OpenStudy (nubeer):

let denominator to be x, numerator = y \[\frac{ y+1 }{ x }= \frac{ 2 }{ 3 }\] Solve this and make an equation. Cnan u do it?

OpenStudy (nubeer):

after u do this for second part it says numerator and denominatoe both increase \[\frac{ y+5 }{ x+5 }= \frac{ 7 }{ 10 }\]

OpenStudy (nubeer):

after u solve this you will get 2 equations.. solve them simultaneously

OpenStudy (anonymous):

First part cannot be done by algebra...

OpenStudy (anonymous):

10/15 = 2/5 Therefore the answer is 9/15

OpenStudy (anonymous):

First is 2x=3y+3

OpenStudy (anonymous):

Second is 7x=10y+15

OpenStudy (anonymous):

@nubeer x=-12 and y=-9

OpenStudy (anonymous):

The fraction is 3/4?

OpenStudy (anonymous):

3/5

OpenStudy (nubeer):

please find the values of x and y again.. i think something wrong there

OpenStudy (anonymous):

After simplifcation

OpenStudy (nubeer):

it says 2/3 not 2/5 @SGREBORN

OpenStudy (anonymous):

Ooops its y=9 and x=15 sorry im wrong in signs

OpenStudy (anonymous):

Sorry my bad..first part we cant use algebra.... instead to use logic

OpenStudy (anonymous):

@nubeer 9/15 or 3/5?

OpenStudy (nubeer):

9/15

OpenStudy (anonymous):

Thanks!!!